# Bias-Corrected Data Synthesis for Imbalanced Learning

### v1

@@ 4.1 Mean Shift Model
Consider the loss function $\ell(\bm{x},y;\bm{\beta})=-y\cdot\bm{x}^{\top}\bm{\beta}+\log(1+\exp(\bm{x}^{\top}\bm{\beta}))$. For the three methods – using raw data, synthetic-augmented data, and synthetic-augmented data with bias correction – we train models by minimizing the respective loss functions given in Equations (2.1), (2.2), and (2.8) for 100 epochs. The entire simulation procedure is repeated over 100 replicates to reduce the impact of randomness. The resulting evaluation metrics for the three methods are summarized in Table 1.
@@
Table 1: Performance metrics (recall, precision, F1-score) evaluated on raw and synthetic-augmented data (with/without bias correction) across varying distributions based on the mean shift model. The results are based on 100 simulations and bold values indicate the top-performing method per metric.
$t(2)$ $\mathcal{N}(\cdot,1)$ $Logistic(\cdot,5)$
Raw SMOTE Bias Corr Raw SMOTE Bias Corr Raw SMOTE Bias Corr
Recall 0.5942 0.7240 0.7317 0.6404 0.7916 0.7962 0.5449 0.5497 0.5553
Precision 0.0059 0.0071 0.0072 0.0065 0.0080 0.0080 0.0055 0.0056 0.0056
$F_{\beta}$ Score 0.5927 0.7221 0.7298 0.6388 0.7896 0.7942 0.5435 0.5484 0.5539
@@
Based on an analysis of 100 simulations across varying data distributions in Table 1, the application of the SMOTE synthetic data generator produced a marked performance improvement over using raw data alone, as measured by recall, precision, and $F_{\beta}$-score. Moreover, incorporating the bias correction technique on top of SMOTE leads to an additional improvement, indicating that the combined approach enhances model robustness and generalization. These results demonstrate that while SMOTE remains an effective foundation for handling data imbalance, integrating bias correction further refines the model’s predictive performance, yielding consistent gains across multiple evaluation metrics.
To further demonstrate the effectiveness of the bias correction method on synthetic data, we conduct simulations using some other synthetic generators that can introduce substantial bias. Although directly using the synthetic data from these generators already improves performance compared with methods relying solely on raw data, applying the bias correction procedure leads to a further and substantial performance gain. The results are summarized in Table 2.
Table 2: Performance metrics (Recall, Precision, F1-score, Jaccard index) evaluated on raw and synthetic data (with/without bias correction) across three synthetic methods (Gaussian mixture, perturbed sampling and biased SMOTE) based on the mean shift model. The results are based on 100 simulations and bold values indicate the top-performing method per metric.
Gaussian-Mixture Perturbed-Sampling Biased-SMOTE
Raw Synthetic Bias Corr Raw Synthetic Bias Corr Raw Synthetic Bias Corr
Recall 0.0968 0.3161 0.3247 0.0980 0.2845 0.3083 0.0940 0.8672 0.9074
Precision 0.4951 0.4982 0.4990 0.5058 0.4969 0.4948 0.4989 0.5031 0.5030
F1 0.1616 0.3782 0.3841 0.1639 0.3501 0.3701 0.1576 0.6363 0.6469
Jaccard 0.0881 0.2367 0.2415 0.0895 0.2158 0.2304 0.0859 0.4670 0.4784
As presented in Table 2, the generation of synthetic data itself provided a significant performance improvement in recall, F1-score, and Jaccard index over models trained solely on raw data, with precision remaining similar. However, in contrast to high-quality generators, these poorer methods introduced substantial bias, which is evidenced by a further marked improvement in performance after the application of a bias correction technique. This demonstrates that for such suboptimal synthetic data, the bias correction procedure is a critically important step that successfully mitigates inherent biases and leads to the best overall model performance.
4.2 Non-linear Classification
To further evaluate the effectiveness and robustness of the proposed bias correction technique under diverse conditions, we conducted a series of controlled simulations using four non-linear classification settings. Each setting represents a distinct geometric relationship between the majority and minority classes, designed to capture a range of challenging data imbalance scenarios. The datasets were constructed with varying degrees of class overlap, non-convexity, and variance, providing a comprehensive test bed for assessing model robustness. For each configuration, synthetic samples were generated using the SMOTE algorithm followed by a bias correction step. The classification performance was evaluated in terms of the $F_{\beta}$ score, averaged over 100 independent runs to ensure statistical reliability.
@@ 4.2 Non-linear Classification
4.3 Sigmoid Bernoulli Model
In this section, a simulation study on the sigmoid Bernoulli model is conducted. We mainly study the behavior for four exponential family distributions (Gaussian distribution, Gumbel distribution, location-scale $t$ distribution and HSD) with a relatively high-quality synthetic generator (SMOTE) and a suboptimal synthetic generator with random sampling with noise (perturbed sampling). Two distributions outside of the exponential family (Laplace distribution and logistic distribution) are also tested on the SMOTE-based synthetic data and give good results, demonstrating the robustness of the proposed method to various distributions. The sample size is set to n=1,000n=1{,}000, and the dimension to $d=10$. The mean squared errors are reported based on 100 simulation runs.
@@ 4.3 Sigmoid Bernoulli Model
Table 4: Estimation error of parameter $\bm{\beta}$ evaluated on raw and synthetic data (with/without bias correction) with two synthetic methods (Perturbed Sampling and Gaussian Mixture) across varying distributions based on sigmoid Bernoulli model. The results are based on 100 simulations and bold values indicate the top-performing method.
Gaussian Mixture Perturbed Sampling
Raw Synthetic Bias Corr Raw Synthetic Bias Corr
Gumbel 2.459 2.418 2.325 2.458 2.438 2.219
HSD 2.511 2.472 2.440 2.512 2.471 2.297
Logistic 2.349 2.507 2.311 2.347 2.658 2.205
The parameter estimation error for $\beta$ in Table 4 demonstrates that the two relatively low-quality synthetic data generators—perturbed sampling and Gaussian mixture—can, in fact, degrade estimation accuracy, yielding higher errors than those obtained using the raw data alone. This deterioration occurs because these generators introduce systematic bias and distort the underlying data distribution, leading to unreliable parameter estimates. However, applying the bias correction technique effectively eliminates this detrimental effect and not only restores performance to the baseline level but also surpasses the accuracy achieved with raw data. This outcome underscores the robustness and corrective strength of the proposed method: even when the synthetic data generator fails to model the true distribution faithfully, bias correction can compensate for these deficiencies and produce more accurate and stable parameter estimates across diverse data settings.
4.4 Average Treatment Effect Estimation
In this section, we conduct a simulation study to evaluate the performance of three methods for estimating the ATE. We compare the standard estimator applied to the raw data with estimators applied to data augmented by SMOTE with and without bias correction. The covariates $\bm{X}$ are generated from four distributions: t(6), t(4), Logistic, and Laplace. The results are summarized in Figure 3.

### v2

@@ B.1 Mean Shift Model
Consider the loss function $\ell(\bm{x},y;\bm{\beta})=-y\cdot\bm{x}^{\top}\bm{\beta}+\log(1+\exp(\bm{x}^{\top}\bm{\beta}))$. For the three methods, using raw data, synthetic-augmented data, and synthetic-augmented data with bias correction, we train models by minimizing the respective loss functions given in Equations (2.1), (2.2), and (2.8) for 100 epochs.
Due to the imbalanced nature of the simulated data, overall accuracy is not an appropriate performance metric and can be misleading, particularly when class proportions are modified through upsampling. Therefore, we evaluate all methods using metrics that are more informative for minority-class performance, including recall, precision, and $F_{\beta}$-score with $\beta=1/\pi_{1}$. The $F_{\beta}$ score is defined as
$F_{\beta}=\frac{(1+\beta^{2}){\rm TP}}{(1+\beta^{2}){\rm TP}+{\rm FP}+\beta^{2}{\rm FN}},$
where ${\rm TP},{\rm FP}$ and $\rm FN$} represent the number of true positives, false positives and false negatives, respectively. In this case we choose $\beta=1/\pi_{1}$, to account for class imbalance, so that false negatives are penalized more heavily when the minority class is rare. Importantly, all upsampling procedures are applied exclusively to the training samples. Performance is evaluated on test sets that preserve the original class distribution, so that the reported results reflect performance under the true data-generating mechanism rather than the resampled training space.
@@
Table 1: Performance metrics (recall, precision, F1-score) evaluated on raw and synthetic-augmented data, with/without bias correction (denoted as SMOTE and BC, respectively), across varying distributions based on the mean shift model. The results are based on 100 simulations and bold values indicate the top-performing method per metric.
$t(2)$ $\mathcal{N}(\cdot,1)$ $Logistic(\cdot,1)$
Recall Precision $F_{\beta}$ Score Recall Precision $F_{\beta}$ Score Recall Precision $F_{\beta}$ Score
Raw 0.5942 0.0059 0.5927 0.6404 0.0065 0.6388 0.6384 0.0065 0.6368
SMOTE 0.7240 0.0071 0.7221 0.7916 0.0080 0.7896 0.7927 0.0081 0.7907
Bootstrap 0.7163 0.0071 0.7145 0.7825 0.0079 0.7806 0.7931 0.0081 0.7912
Reweight 0.6703 0.0066 0.6686 0.7271 0.0073 0.7253 0.7196 0.0073 0.7178
SMOTE-BC 0.7317 0.0072 0.7298 0.7962 0.0080 0.7942 0.7972 0.0081 0.7952
@@
The resulting evaluation metrics for the three methods are summarized in Table 1. Based on an analysis of 100 simulations across varying data distributions in Table 1, the application of the SMOTE synthetic data generator produced a marked performance improvement over using raw data alone, as measured by recall, precision, and $F_{\beta}$-score. Moreover, incorporating the bias correction technique on top of SMOTE leads to an additional improvement, indicating that the combined approach enhances model robustness and generalization. These results demonstrate that while SMOTE remains an effective foundation for handling data imbalance, integrating bias correction further refines the model’s predictive performance, yielding consistent gains across multiple evaluation metrics.
B.2 Non-linear Classification
To further evaluate the effectiveness and robustness of the proposed bias correction technique under diverse non-linear decision boundaries, we consider four controlled two-dimensional classification settings, each representing a distinct geometric relationship between the majority and minority classes as illustrated in Figure 3. In each replicate, we generate $n=1,000$ observations with $d=2$ features and labels $Y\in\{0,1\}$ sampled i.i.d. from $\mathrm{Bernoulli}(p_{1})$ with minority prevalence $\pi_{1}=0.05$. The four configurations vary in class overlap, non-convexity, and variance structure: one setting produces two elongated clouds with opposing correlation directions, yielding intersecting diagonal structures, another places the majority class on a noisy ring surrounding a minority strip, a third yields overlapping clusters with unequal spread and a location shift, and a fourth forms a clearly non-convex majority distribution with two lobes while the minority class occupies a horizontal band. For evaluation, we generate an independent test set of size $0.5n$ with balanced labels $\mathrm{Bernoulli}(0.5)$ to assess performance without class-prior effects. For each configuration, we generate synthetic samples using SMOTE with $k=5$ nearest neighbors and $n_{s}=\lfloor(n_{0}-n_{1})\cdot\texttt{syn\_rate}\rfloor$ (with $\texttt{syn\_rate}=0.5$), and then compare standard synthetic augmentation against the bias-corrected objective. We train the classifier using a two-hidden-layer ReLU network with 12 units per layer, optimized by SGD for 500 epochs (learning rate $0.05$ and weight decay $10^{-4}$). Performance is summarized by the $F_{\beta}$ score with $\beta=1/\pi_{1}$, averaged over 100 independent runs.
For each configuration, we compare standard synthetic augmentation with the bias-corrected objective using the same classifier. Performance is evaluated on an independent balanced test set using the $F_{\beta}$ score, with results averaged over $100$ replicates.
@@ B.2 Non-linear Classification
B.3 Sigmoid Bernoulli Model
In this section, a simulation study on the sigmoid Bernoulli model is conducted to further evaluate the effectiveness and robustness of the proposed bias correction technique under a broad range of data-generating distributions and synthetic generators. Each replicate uses sample size n=1,000n=1{,}000 and dimension $d=10$. Covariates $\bm{X}\in\mathbb{R}^{d}$ are generated from one of several parametric families with a location vector $\mu$ and a common scale factor, and responses follow a sigmoid Bernoulli model
$\mathbb{P}(Y=1\mid\bm{X})=\sigma(\bm{X}^{\top}\bm{\beta}^{\star}),\qquad\bm{\beta}^{\star}=-\mathbf{1}_{d},$
where $\sigma(\cdot)$ denotes the logistic sigmoid function. We mainly study four exponential-family settings for elements of $\bm{X}$: Gaussian, Gumbel, location-scale $t$ (with $t_{4}$ noise), and HSD, and we additionally test two distributions outside the exponential family (Laplace and logistic) to assess robustness. For each distribution, $\bm{\mu}$ is chosen to induce a highly imbalanced label distribution (with a minority rate around $\pi_{1}=0.05$), and the data are split into training/validation/test sets with proportions $60\%/20\%/20\%$. To address imbalance, use SMOTE with $k=5$ nearest neighbors and compare it with the proposed bias correction method. The number of synthetic samples is set to $\tilde{n}=\lfloor(n_{0}-n_{1})\cdot\texttt{syn\_rate}\rfloor$ with $\texttt{syn\_rate}=10\pi_{1}(1-2\pi_{1})$. Performance is evaluated by the estimation error of $\bm{\beta}$, averaged over $100$ independent simulation runs.
@@ B.3 Sigmoid Bernoulli Model
B.4 Average Treatment Effect Estimation
In this section, we conduct a simulation study to evaluate the performance of three methods for estimating the average treatment effect (ATE) under covariate imbalance and distributional heterogeneity. For each run, we generate $n=2000$ observations with $d=2$ covariates $\bm{X}$, and assign a binary treatment $Z$ according to a logistic propensity model
$\mathbb{P}(Z=1\mid\bm{X})=\sigma(\beta_{z,0}+\bm{\beta}_{z}^{\top}\bm{X}),$
with $\beta_{z,0}=-1.6$ and $\bm{\beta}_{z}=\mathbf{1}_{d}$, which induces an imbalanced treated proportion. Potential outcomes follow linear models,
$Y(1)=a_{1}+\bm{X}^{\top}\bm{\mu}_{1}+\varepsilon_{1},\qquad Y(0)=a_{0}+\bm{X}^{\top}\bm{\mu}_{0}+\varepsilon_{0},$
with $a_{1}=a_{0}=1$, $\bm{\mu}_{1}=(2,1)^{\top}$, $\bm{\mu}_{0}=\bm{\mu}_{1}$, and independent Gaussian noise with standard deviation $\sigma_{y}=1$. The observed outcome is Y=ZY(1)+(1Z)Y(0)Y=ZY(1)+(1-Z)Y(0), and the true ATE is $\tau=\mathbb{E}[Y(1)-Y(0)]$. We compare (i) the standard AIPW estimator computed on the raw data, (ii) an AIPW estimator that fits the propensity model after augmenting the treated group with SMOTE-generated synthetic covariates, and (iii) the same SMOTE augmentation combined with the proposed bias-corrected propensity fitting. The number of synthetic treated samples is chosen to target a treated share of $0.5$, and SMOTE uses $k=5$ nearest neighbors. Covariates $X$ are generated from four distributions: $t(4)$, $t(6)$, logistic, and Laplace. Performance is summarized by the RMSE of $\widehat{\tau}$ over $100$ independent runs, and the results are reported in Figure 4.